# Non-Markovian character and irreversibility of real-time quantum many-body dynamics

### v1

@@ Abstract
The presence of pairing correlations within the time-dependent density functional theory (TDDFT) extension to superfluid systems, is tantamount to the presence of a quantum collision integral in the evolution equations, which leads to an obviously non-Markovian behavior of the single-particle occupation probabilities, unexpected in a traditional quantum extension of kinetic equations. The quantum generalization of the Boltzmann equation, based on a collision integral in terms of phase-space occupation probabilities, is the most used approach to describe nuclear dynamics and which by construction has a Markovian character. By contrast, the extension of TDDFT to superfluid systems has similarities with the Baym and Kadanoff kinetic formalism, which however is formulated with much more complicated evolution equations with long-time memory terms and non-local interactions.
The irreversibility of quantum dynamics is properly characterized using the canonical wave functions/natural orbitals, and the associated canonical occupation probabilities, which provide the smallest possible representation of any fermionic many-body wave function. In this basis, one can evaluate the orbital entanglement entropy, which is an excellent measure of the non-equilibrium dynamics of an isolated system.
To explore the phenomena of memory effects and irreversibility, we investigate the use of canonical wave functions/natural orbitals in nuclear many-body calculations, assessing their utility for static calculations, dynamics, and symmetry restoration. As the number of single-particle states is generally quite large, it is highly desirable to work in the canonical basis whenever possible, preferably with a cutoff. We show that truncating the number of canonical wave functions can be a valid approach in the case of static calculations, but that such a truncation is not valid for time-dependent calculations, as it leads to the violation of continuity equation, energy conservation, and other observables, as well as an inaccurate representation of the dynamics. Indeed, in order to describe the dynamics of a fissioning system within a TDDFT framework all the canonical states must be included in the calculation. Finally, we demonstrate that the canonical representation provides a very efficient basis for performing symmetry projections.
@@ I Introduction
In Section II we will describe how canonical wave functions/natural orbitals are defined, subsection II.1, and relevant aspects of the Bloch-Messiah decomposition, subsection II.2. In Section III we will discuss generalized Bogoliubov quasiparticles, which become relevant in reactions, when both partners are superfluid, as in the case of nuclear fission or collision between heavy-ions. In Section IV we will discuss several particular aspects which are relevant in the subsequent analysis of the time-dependent equations for fermionic superfluids. In Section V we describe some new aspects of the particle projection for fermionic superfluids, which will be illustrated in the case of nuclear fission in the following section. In subsection VI.1 we discuss under what circumstances the use of a reduced set of canonical wave functions is beneficial. In subsection VI.2, we discuss the non-Markovian character of the fission dynamics. In subsection VI.3, we discuss the relevance of particle projection and also the use of the reduced set of canonical wave functions in dynamic simulations. And in subsection VI.4 we will illustrate the irreversible fission dynamics and its characterization by the means of the orbital entanglement entropy. In Section VII we will summarize our main findings.
@@ II Canonical wave functions
The set of quasi-particle wave functions is twice the size of the set of canonical wave functions . These two sets of quasi-particle wave functions however are related and one can derive one set from the other and vice versa, see Ref. Bulgac et al. (2023) and below. Practice shows that using a significantly reduced set of functions , with occupation numbers is often sufficient to represent the many-body wave function with sufficient accuracy Bulgac et al. (2023). The quasi-particle wave functions are very useful to describe low-energy excitations of the many-body system and their corresponding eigenvalues play a similar role as the particle eigenstates in a normal system.
@@ II.1 Relations between the canonical wave functions and the quasi-particle wave functions
The time-reversal symmetry between the two of the doublet is not a necessary condition in order to uniquely evaluate the anomalous density, as one can show that a unitary transformation between the two time-reversed canonical wave functions do not change the Cooper pair wave function or the anomalous density
@@ III Generalized Bogoliubov quasi-particles
And this “communication” problem and exchange of “information” between spatially separated superfluid systems is even more complicated than the mere influence of the pairing field of one system on another system. For the sake of the argument imagine that two hypothetical nuclei with zero electric charge (in order to exclude long-range Coulomb interaction) or two fermionic neutral superfluid cold atomic clouds are separated by a distance much larger than the average interparticle separation in each system or larger than the range of the interparticle interaction. Such systems are still able to “communicate” due to the fact that the -quasi-particle wave function components are continuum wave functions. Thus any change in one system, due to its own quantum evolution due to the short-range interaction between the particles localized in that system, is “communicated” via the -quasi-particle wave function components to the other system, which in principle could be at the other end of the universe. One might argue that this is simply an artifact of mixing systems with different particle numbers in the HFB approximation. However, this argument cannot be valid for systems with a finite number of particles, where the probability to have a system with a very large number of particle is exponentially small and insufficient to bring in material contact the two systems, as can be easily shown by performing a particle projection, see Section V.
Moreover, even after particle projection of the HFB equations the anomalous densities and the corresponding pairing fields have tails much longer than the matter distribution of the two subsystems. It is not clear yet whether within a theoretical treatment of the pairing correlations, where the particle numbers are exactly conserved, the pairing field will cease to have longer tails than the regular mean field. Even after particle projection the tails of the pairing fields are not affected, see Section V. Using the two type of generalized Bogoliubov quasi-particle described above one can address these questions at least within the HFB approximation after particle projection.
@@ IV Time-dependent equations
The main difference with the static case is that if one starts with a system of canonical quasi-particle states at any time , at the next time step the new set of quasi-particle states ceases to be canonical. In practice this is not a problem if one uses at all times the full set of quasi-particle states. However, it is known that if at any given time one chooses to represent the density matrices using canonical quasi-particle states, for a sufficient accuracy one can obtain their representation with a significantly reduced number of states. For example, in performing fission dynamics simulations of an actinide nucleus one needs to represent the quasi-particle states on a spatial lattice of typical size Bulgac et al. (2019, 2020); Jin et al. (202), in which case the total number of proton and neutron quasi-particle states is . At any time during the time evolution one can however introduce the set of canonical quasi-particle states and represent the same normal and anomalous densities with a comparable numerical accuracy with about 5,000 states or less, see Ref. Bulgac et al. (2023) and also Section VI. As a matter of fact, in static calculations accurate solutions were obtained with a significantly reduced set of canonical wave functions Chen et al. (2022), for example in the case of Pu these authors reproduced the static states binding energies with at most 400 neutron orbitals and 300 proton orbitals, but see also Section VI.
@@ V Particle projection
where , which in terms of canonical QPWFs
@@
The particle distribution is given by where the particle projector is
One can show that the -particle projected many-body wave function is given by
@@
is the wave function of a “Cooper pair,” also known in the literature as a geminal. The number projected wave function is a sum of linearly independent -particle Slater determinants.
For an arbitrary operator the particle projected value in a state with exactly particles is defined as Bulgac (2021)
where the fact that the operator is a projector was used. The particle projected number density matrix is
@@
In order to evaluate various particle projected number densities it is helpful to introduce an additional new set of particle projected occupation probabilities
with and . The particle projected density matrices have then the form
@@
For Eqs. (142-144) lead to the corresponding particle unprojected number density matrices. In Eqs. (142-144) the probabilities are evaluated as in Eq. (130), but with a missing contribution from quasi-particle state
From these expressions one can write in straightforward manner the corresponding expressions for the particle projected normal number, kinetic, current, spin, spin-current and anomalous densities. What is notable about Eqs. (142-144) is that the contribution of the troublesome self-interacting term is excluded in all these one-body densities and their use is free of the self-energy problem widely discussed in literature Stringari and Brink (1978); Perdew and Zunger (1981); Stoitsov et al. (2007); Bender et al. (2003); Dobaczewski et al. (2007); Duguet et al. (2009); Bender et al. (2009); Lacroix et al. (2009); Hupin et al. (2011); Duguet et al. (2015). Hupin et al. (2011) have obtained similar, though somewhat more complicated, relations.
@@
Therefore, the only source for a self-interacting energy contribution can appear from the parameterization of the exchange contribution in DFT, a problem which is absent when using the particle projected number density. In the treatment of the unitary Fermi gas in Refs. Bulgac (2007); Bulgac et al. (2012) a relatively large discrepancy was observed when comparing the Quantum Monte Carlo (QMC) results to the SLDA results only in the case of two fermions with opposite spins, where exchange is absent, but the self-interacting energy is present in the energy density functional, as the unprojected particle number density was used. For larger particle numbers the agreement between QMC and SLDA results was always within the QMC statistical errors.
With the particle projected densities one can evaluate the energy of a nucleus for a fixed particle number within DFT, as in the ideology of DFT the energy is well defined once the energy density functional is known and the relevant number densities as well. In the mean field framework, theorists refer to this recipe as the projection either before or after energy variation. The difference between “variation before” or “variation after” recipes is only in the manner the quasi-particle wave functions are determined. In the “after variation” approach, one solves the unprojected mean field equations and determines the energy minimum and the particle and/or the angular and/or parity projection are performed after the minimum is found and the projected energy is calculated. In the projection-before-variation approach, one performs the variation on the projected energy functional at first and determines the minimum energy subsequently. In determining the next step in a self-consistent procedure one can use however the number projected densities instead of the particle unprojected energies, and in principle arrive at the same result. We want to stress that particle projection is very inexpensive to perform and as a result one can perform DFT calculation for superfluid systems with exact particle numbers and in this case the self-energy terms are entirely absent. In “parameterizing” the energy density function therefore one can treat independently the exchange and the pairing correlation energies, thus removing a difficulty with the theoretical treatment of such systems Sheikh et al. (2021).
@@ VI.1 On the use of a reduced set of canonical wave functions
All our fission dynamics simulations were perform using a fully self-consistent set of quasi-particle wave functions on a spatial lattice fm with a lattice constant of fm, corresponding to a momentum cutoff in one direction MeV/c, corresponding to a nucleon kinetic energy of more 180 MeV (or more than 360 MeV kinetic energy in the center of mass for two colliding nucleons), significantly larger than the Fermi energy of symmetric nuclear matter, and very close to the upper energy considered in chiral EFT approaches. The self-consistent equations were solved for the nuclear density functional SeaLL1 Bulgac et al. (2018), at first using the HFBTHO code Navarro Perez et al. (2017), after which we ported the proton and neutron number, current, kinetic energy and anomalous densities to our 3D spatial lattice and continued the self-consistent procedure until full convergence. The size of the HFBTHO basis set for either proton or neutron system was about 8,000 compared to the basis set of our 3D spatial lattice, which was 108,000. In Table 1 we display some properties of the initial state at the top of the outer fission barrier, evaluated either with the full set of quasi-particle wave functions obtained in the self-consistent procedure described here, or using the entire set of canonical wave functions instead, or only a reduced set of canonical wave functions, selected with the largest occupation probabilities. As expected the last two entries in the Table 1 are identical. It is important to compare however the initial converged energy obtained with the HFBTHO code MeV, which is considerably above the energy MeV we obtain on the 3D lattice.
# States | Energy | N | Z | ||
@@
1,000 | -1780.64 | 143.97 | 92.00 | 169.76 | 22.29 |
2,000 | -1782.64 | 143.99 | 92.00 | 169.78 | 22.29 |
5,000 | -1785.31 | 144.01 | 92.00 | 169.79 | 22.30 |
50,000 | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
108,000 (Canonical) | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
108,000 (Standard) | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
# States | ||||||
@@
1000 | 91.98 | 91.64 | 0.34 | 144.00 | 139.92 | 4.07 |
2000 | 91.99 | 91.66 | 0.33 | 144.02 | 140.06 | 3.96 |
5000 | 91.99 | 91.69 | 0.30 | 144.03 | 140.33 | 3.70 |
50,000 | 91.99 | 91.97 | 0.017 | 144.03 | 143.23 | 0.80 |
108,000 | 91.99 | 91.99 | 144.03 | 144.03 |
On the other hand the total initial energy, particle numbers, and deformations are reproduced adequately only when at least 5,000 canonical wave functions are taken into account. One might have naively assumed that using 500 canonical wave functions for both proton and neutron systems the properties of the initial state would be quite well reproduced, as Chen et al. (2022) expected. These authors used used 400 and 300 canonical wave functions for the neutron and proton systems in order to evaluate the energy of the fission isomer Pu, which should be compared with our claim that at least 5,000 canonical wave functions are needed to obtain a comparable accuracy for the total energy of a similar nucleus. Reducing the number of canonical wave functions from 5,000 to 2,000, see Table 1, leads to a significant error in the energy. In Table 2 we have evaluated the final proton and neutron numbers, where we used the entire set of canonical wave functions to evolve the system, see penultimate row in Table 1, within a reduced set of quasi-particle wave functions with the corresponding largest initial occupation probabilities, in the range from 500 to 50,000, for both neutron and proton subsystems. This demonstrates that the naive choice of a reduced set initial quasi-particle wave functions, which however reproduced accurately the initial neutron and proton numbers cannot reproduce the corresponding final total neutron and proton numbers.
@@
A feature we generally observe in our time-dependent simulations is that the quality of the initial state plays a significant role in the accuracy of the entire simulation until the FFs are fully separated. Better converged solutions for the initial state lead to significant increase in the accuracy of the simulations, judged by either the conservation of the particle numbers or total energy. Comparing the results obtained with either the full set of quasi-particle wave functions, last row in Table 1, with the full set of the canonical wave functions we observed relative errors at the level of , after performing about 30,000 time-steps, depending on the quantity. The total particle number is conserved at the level of and the total energy at the level of noticeably less than 1 KeV, thus a relative error less than .

### v2

@@ Abstract
The presence of pairing correlations within the time-dependent density functional theory (TDDFT) extension to superfluid systems, is tantamount to the presence of a quantum collision integral in the evolution equations, which leads to an obviously non-Markovian behavior of the single-particle occupation probabilities, unexpected in a traditional quantum extension of kinetic equations. The quantum generalization of the Boltzmann equation, based on a collision integral in terms of phase-space occupation probabilities, is the most used approach to describe nuclear dynamics and which by construction has a Markovian character. By contrast, the extension of TDDFT to superfluid systems has similarities with the Baym and Kadanoff kinetic formalism, which however is formulated with much more complicated evolution equations with long-time memory terms and non-local interactions. The irreversibility of quantum dynamics is properly characterized using the canonical wave functions/natural orbitals, and the associated canonical occupation probabilities, which provide the smallest possible representation of any fermionic many-body wave function. In this basis, one can evaluate the orbital entanglement entropy, which is an excellent measure of the non-equilibrium dynamics of an isolated system. To explore the phenomena of memory effects and irreversibility, we investigate the use of canonical wave functions/natural orbitals in nuclear many-body calculations, assessing their utility for static calculations, dynamics, and symmetry restoration. As the number of single-particle states is generally quite large, it is highly desirable to work in the canonical basis whenever possible, preferably with a cutoff. We show that truncating the number of canonical wave functions can be a valid approach in the case of static calculations, but that such a truncation is not valid for time-dependent calculations, as it leads to the violation of continuity equation, energy conservation, and other observables, as well as an inaccurate representation of the dynamics. Indeed, in order to describe the dynamics of a fissioning system within a TDDFT framework all the canonical states must be included in the calculation. Finally, we demonstrate that the canonical representation provides a very efficient basis for performing symmetry projections.
@@ I Introduction
In Section II we will describe how canonical wave functions/natural orbitals are defined, subsection II.1, and relevant aspects of the Bloch-Messiah decomposition, subsection II.2. In Section III we will discuss generalized Bogoliubov quasiparticles, which become relevant in reactions, when both partners are superfluid, as in the case of nuclear fission or collision between heavy-ions. In Section IV we will discuss several particular aspects which are relevant in the subsequent analysis of the time-dependent equations for fermionic superfluids. In Section V we describe some new aspects of the particle number projection for fermionic superfluids, which will be illustrated in the case of nuclear fission in the following section. In subsection VI.1 we discuss under what circumstances the use of a reduced set of canonical wave functions is beneficial. In subsection VI.2, we discuss the non-Markovian character of the fission dynamics. In subsection VI.3, we discuss the relevance of particle number projection and also the use of the reduced set of canonical wave functions in dynamic simulations. And in subsection VI.4 we will illustrate the irreversible fission dynamics and its characterization by the means of the orbital entanglement entropy. In Section VII we will summarize our main findings.
@@ II Canonical wave functions
The set of quasi-particle wave functions is twice the size of the set of canonical wave functions . These two sets of quasi-particle wave functions however are related and one can derive one set from the other and vice versa, see Ref. Bulgac et al. (2023) and the text below. Practice shows that using a significantly reduced set of functions , with occupation numbers is often sufficient to represent the many-body wave function with sufficient accuracy Bulgac et al. (2023). The quasi-particle wave functions are very useful to describe low-energy excitations of the many-body system and their corresponding eigenvalues play a similar role as the particle eigenstates in a normal system.
@@ II.1 Relations between the canonical wave functions and the quasi-particle wave functions
The time-reversal symmetry between the two of the doublet is not a necessary condition in order to uniquely evaluate the anomalous density, as one can show that a unitary transformation between the two time-reversed canonical wave functions does not change the Cooper pair wave function or the anomalous density
@@ III Generalized Bogoliubov quasi-particles
And this “communication” problem and exchange of “information” between spatially separated superfluid systems is even more complicated than the mere influence of the pairing field of one system on another system. For the sake of the argument imagine that two hypothetical nuclei with zero electric charge (in order to exclude long-range Coulomb interaction) or two fermionic neutral superfluid cold atomic clouds are separated by a distance much larger than the average interparticle separation in each system or larger than the range of the interparticle interaction. Such systems are still able to “communicate” due to the fact that the -quasi-particle wave function components are continuum wave functions. Thus any change in one system, due to its own quantum evolution due to the short-range interaction between the particles localized in that system, is “communicated” via the -quasi-particle wave function components to the other system, which in principle could be at the other end of the universe. One might argue that this is simply an artifact of mixing systems with different particle numbers in the HFB approximation. However, this argument cannot be valid for systems with a finite number of particles, where the probability to have a system with a very large number of particle is exponentially small and insufficient to bring in material contact the two systems, as can be easily shown by performing a particle number projection, see Section V.
Moreover, even after particle number projection of the HFB equations the anomalous densities and the corresponding pairing fields have tails much longer than the matter distribution of the two subsystems. It is not clear yet whether within a theoretical treatment of the pairing correlations, where the particle numbers are exactly conserved, the pairing field will cease to have longer tails than the regular mean field. Even after particle number projection the tails of the pairing fields are not affected, see Section V. Using the two type of generalized Bogoliubov quasi-particle described above one can address these questions at least within the HFB approximation after particle number projection.
@@ IV Time-dependent equations
The main difference with the static case is that if one starts with a system of canonical quasi-particle states at any time , at the next time step the new set of quasi-particle states ceases to be canonical. In practice this is not a problem if one uses at all times the full set of quasi-particle states. However, it is known that if at any given time one chooses to represent the density matrices using canonical quasi-particle states, for a sufficient accuracy one can obtain their representation with a significantly reduced number of states. For example, in performing fission dynamics simulations of an actinide nucleus one needs to represent the quasi-particle states on a spatial lattice of typical size Bulgac et al. (2019, 2020); Jin et al. (202), in which case the total number of proton and neutron quasi-particle states is . At any time during the time evolution one can however introduce the set of canonical quasi-particle states and represent the same normal and anomalous densities with a comparable numerical accuracy with about 5 000 states or less, see Ref. Bulgac et al. (2023) and also Section VI. As a matter of fact, in static calculations accurate solutions were obtained with a significantly reduced set of canonical wave functions Chen et al. (2022), for example in the case of 240Pu these authors reproduced the static states binding energies with at most 400 neutron orbitals and 300 proton orbitals, but see also Section VI.
@@ V Particle-number projection
where , which in terms of canonical quasi-particle wave functions (QPWFs)
@@
where is a gauge angle. The particle number distribution is given by where the particle projector is
One can show that the -particle number-projected many-body wave function is given by
@@
is the wave function of a “Cooper pair,” also known in the literature as a geminal. The number projected wave function is a sum of linearly independent -particle Slater determinants. is a state of “Cooper pairs” with proper normalization.
For an arbitrary operator the particle number projected value in a state with exactly particles is defined as Bulgac (2021)
where the fact that the operator is a projector was used. The particle number projected number density matrix is
@@
In order to evaluate various particle number projected number densities it is helpful to introduce an additional new set of particle number projected occupation probabilities
with and . The particle number projected density matrices have then the form
@@
For Eqs. (91-93) lead to the corresponding particle number unprojected number density matrices. In Eqs. (91-93) the probabilities are evaluated as in Eq. (79), but with a missing contribution from quasi-particle state
From these expressions one can write in straightforward manner the corresponding expressions for the particle number projected normal number, kinetic, current, spin, spin-current and anomalous densities. What is notable about Eqs. (91-93) is that the contribution of the troublesome self-interacting term is excluded in all these one-body densities and their use is free of the self-energy problem widely discussed in literature Stringari and Brink (1978); Perdew and Zunger (1981); Stoitsov et al. (2007); Bender et al. (2003); Dobaczewski et al. (2007); Duguet et al. (2009); Bender et al. (2009); Lacroix et al. (2009); Hupin et al. (2011); Duguet et al. (2015). Hupin et al. (2011) have obtained similar, though somewhat more complicated, relations.
@@
Therefore, the only source for a self-interacting energy contribution can appear from the parameterization of the exchange contribution in DFT, a problem which is absent when using the particle number projected number density. In the treatment of the unitary Fermi gas in Refs. Bulgac (2007); Bulgac et al. (2012) a relatively large discrepancy was observed when comparing the Quantum Monte Carlo (QMC) results to the SLDA results only in the case of two fermions with opposite spins, where exchange is absent, but the self-interacting energy is present in the energy density functional, as the unprojected particle number density was used. For larger particle numbers the agreement between QMC and SLDA results was always within the QMC statistical errors.
With the particle number projected densities one can evaluate the energy of a nucleus for a fixed particle number within DFT, as in the ideology of DFT the energy is well defined once the energy density functional is known and the relevant number densities as well. In the mean field framework, theorists refer to this recipe as the projection either before or after energy variation. The difference between “variation before” or “variation after” recipes is only in the manner the quasi-particle wave functions are determined. In the “after variation” approach, one solves the unprojected mean field equations and determines the energy minimum and the particle and/or the angular and/or parity projection are performed after the minimum is found and the projected energy is calculated. In the projection-before-variation approach, one performs the variation on the projected energy functional at first and determines the minimum energy subsequently. In determining the next step in a self-consistent procedure one can use however the number projected densities instead of the particle number unprojected energies, and in principle arrive at the same result. We want to stress that particle number projection is very inexpensive to perform and as a result one can perform DFT calculation for superfluid systems with exact particle numbers and in this case the self-energy terms are entirely absent. In “parameterizing” the energy density function therefore one can treat independently the exchange and the pairing correlation energies, thus removing a difficulty with the theoretical treatment of such systems Sheikh et al. (2021).
@@ VI.1 On the use of a reduced set of canonical wave functions
All our fission dynamics simulations were perform using a fully self-consistent set of quasi-particle wave functions on a spatial lattice fm3 with a lattice constant of fm, corresponding to a momentum cutoff in one direction MeV/c, corresponding to a nucleon kinetic energy of more 180 MeV (or more than 360 MeV kinetic energy in the center of mass for two colliding nucleons), significantly larger than the Fermi energy of symmetric nuclear matter, and very close to the upper energy considered in chiral EFT approaches. The self-consistent equations were solved for the nuclear density functional SeaLL1 Bulgac et al. (2018), at first using the HFBTHO code Navarro Perez et al. (2017), after which we ported the proton and neutron number, current, kinetic energy and anomalous densities to our 3D spatial lattice and continued the self-consistent procedure until full convergence. The size of the HFBTHO basis set for either proton or neutron system was about 8 000 compared to the basis set of our 3D spatial lattice, which was 108 000. In Table 1 we display some properties of the initial state at the top of the outer fission barrier, evaluated either with the full set of quasi-particle wave functions obtained in the self-consistent procedure described here, or using the entire set of canonical wave functions instead, or only a reduced set of canonical wave functions, selected with the largest occupation probabilities. As expected the last two entries in the Table 1 are identical. It is important to compare however the initial converged energy obtained with the HFBTHO code MeV, which is considerably above the energy MeV we obtain on the 3D lattice.
No. States | Energy | N | Z | ||
@@
1 000 | -1780.64 | 143.97 | 92.00 | 169.76 | 22.29 |
2 000 | -1782.64 | 143.99 | 92.00 | 169.78 | 22.29 |
5 000 | -1785.31 | 144.01 | 92.00 | 169.79 | 22.30 |
50 000 | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
108 000 (Canonical) | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
108 000 (Standard) | -1785.41 | 144.01 | 92.00 | 169.79 | 22.30 |
No. States | ||||||
@@
1 000 | 91.98 | 91.64 | 0.34 | 144.00 | 139.92 | 4.07 |
2 000 | 91.99 | 91.66 | 0.33 | 144.02 | 140.06 | 3.96 |
5 000 | 91.99 | 91.69 | 0.30 | 144.03 | 140.33 | 3.70 |
50 000 | 91.99 | 91.97 | 0.017 | 144.03 | 143.23 | 0.80 |
108 000 | 91.99 | 91.99 | 144.03 | 144.03 |
On the other hand the total initial energy, particle numbers, and deformations are reproduced adequately only when at least 5 000 canonical wave functions are taken into account. One might have naively assumed that using 500 canonical wave functions for both proton and neutron systems the properties of the initial state would be quite well reproduced, as Chen et al. (2022) expected. These authors used used 400 and 300 canonical wave functions for the neutron and proton systems in order to evaluate the energy of the fission isomer 240Pu, which should be compared with our claim that at least 5 000 canonical wave functions are needed to obtain a comparable accuracy for the total energy of a similar nucleus. Reducing the number of canonical wave functions from 5 000 to 2 000, see Table 1, leads to a significant error in the energy. In Table 2 we have evaluated the final proton and neutron numbers, where we used the entire set of canonical wave functions to evolve the system, see penultimate row in Table 1, within a reduced set of quasi-particle wave functions with the corresponding largest initial occupation probabilities, in the range from 500 to 50 000, for both neutron and proton subsystems. This demonstrates that the naive choice of a reduced set initial quasi-particle wave functions, which however reproduced accurately the initial neutron and proton numbers cannot reproduce the corresponding final total neutron and proton numbers.
@@
A feature we generally observe in our time-dependent simulations is that the quality of the initial state plays a significant role in the accuracy of the entire simulation until the fission fragments (FFs) are fully separated. Better converged solutions for the initial state lead to significant increase in the accuracy of the simulations, judged by either the conservation of the particle numbers or total energy. Comparing the results obtained with either the full set of quasi-particle wave functions, last row in Table 1, with the full set of the canonical wave functions we observed relative errors at the level of , after performing about time-steps, depending on the quantity. The total particle number is conserved at the level of and the total energy at the level of noticeably less than 1 KeV, thus a relative error less than .